# Breaking the Adversarial Robustness-Performance Trade-off in Text Classification via Manifold Purification

### v1

@@ Introduction
The remarkable success of Pre-trained Language Models (PLMs) like BERT across numerous Natural Language Processing (NLP) tasks (devlin2019bert) is shadowed by their pronounced vulnerability to adversarial attacks (jin2020bert; ren2019generating). These attacks introduce minor, often semantically imperceptible, perturbations to input text, yet can trigger catastrophic prediction errors in the model, which becomes especially critical in text classification (TC). (vazquez2024survey; goyal2023survey; li2020bert; wang2021measure) In response, a wide array of defense mechanisms have been developed to enhance model robustness (gao2023dsrm; zheng2024subspace; asl2024robustsentembed; zhu2020freelb). However, their success is often marred by a critical challenge: improving resilience against adversarial inputs frequently comes at the cost of a significant drop in performance on clean, unperturbed data (zhang2020revisiting; wang2021measure). This robustness-accuracy challenge not only poses a significant barrier to the reliable deployment of PLMs in safety-critical applications but also highlights a fundamental gap in our understanding of their internal workings.
@@
Figure 1: Figure 1: An overview of our proposed Manifold-Correcting Causal Flow ($MC^{2}F$) framework. The system first employs a Stratified Riemannian Continuous Normalizing Flow (SR-CNF) to model the manifold of clean data (colored points). It then identifies out-of-distribution (OOD) or adversarial samples (red points) that lie off the manifold. The Purification Solver module subsequently corrects these OOD embeddings by projecting them back onto the clean manifold along the shortest path (geodesic), a process visualized by the black arrows. In-distribution embeddings are passed through without modification.
@@
To overcome this challenge, we move beyond heuristic defenses and investigate the geometric properties of the model’s embedding space. We begin with a critical empirical observation: in the high-dimensional embedding space of a BERT-style encoder, representations of clean text are geometrically separable from their adversarial counterparts generated by prevalent attacks (e.g., TextFooler, BERT-Attack) (subhash2023universal; reif2019visualizing). This suggests that adversarial perturbations induce a measurable geometric and distributional shift, rather than creating indistinguishable overlaps. Consequently, the defense problem can be reframed: instead of a brute-force training task, it becomes a more nuanced geometric challenge of identifying and correcting embeddings that have been perturbed off the ”clean” data distribution.
@@
Building on this insight and inspired by the manifold hypothesis—which posits that high-dimensional data lies on a lower-dimensional manifold—we introduce the Manifold-Correcting Causal Flow ($MC^{2}F$). It is a two-module defense system that operates directly on sentence embeddings, embodying a ”detect-and-correct” method to purify adversarial inputs  (li2023text; moraffah2024adversarial).
@@
Figure 2: 2D visualizations of clean (circles) and attacked (crosses) SST-2 test set embeddings from the BERT encoder using (a) PCA, (b) t-SNE, and (c) UMAP. Across all three dimensionality reduction techniques, the adversarial embeddings form clusters that are visibly distinct from the clean embeddings, suggesting geometric separability. This separation is particularly pronounced in the manifold-aware UMAP projection.
@@
Our work is positioned at the intersection of adversarial defense in NLP and the geometric analysis of embedding spaces. We review the most relevant literature in these areas.
@@
The vulnerability of PLMs has been extensively demonstrated by a variety of adversarial attack methods. These methods typically generate perturbations by replacing, inserting, or deleting characters or words under certain constraints to maintain semantic similarity and fluency. Prominent word-level attack algorithms include TextFooler (jin2020bert), BAE (garg2020bae), and TextBugger (li2018textbugger), which greedily replace words with synonyms that maximize the victim model’s prediction error. More sophisticated methods, such as BERT-Attack (li2020bert), leverage the victim model itself to find more effective substitutes. These attacks serve as standard benchmarks for evaluating the robustness of defense systems, including our own.
@@
Research into adversarial defense has yielded diverse strategies. The predominant paradigm, adversarial training (AT), enhances robustness by augmenting training data with adversarial examples (zhu2020freelb; madry2018towards). However, AT is computationally expensive and notoriously suffers from a trade-off between robust and clean accuracy (zhang2020revisiting; wang2021measure). While recent methods attempt to mitigate this tension (gao2023dsrm; wang2024new), a more fundamental ”illusion of robustness” has been identified, where apparent gains are merely artifacts of gradient masking that can be circumvented by adaptive attacks (raina2024extreme). To avoid these pitfalls, another line of research focuses on input purification, which denoises inputs before classification. This approach has evolved from operating on discrete tokens to purifying continuous embedding representations, for instance, by projecting embeddings onto a learned ”clean” subspace (zheng2024subspace) or using contrastive learning (asl2024robustsentembed). Notably, methods like DAD have demonstrated success in improving both clean and robust accuracy by using Maximum Mean Discrepancy (MMD) for detection and a trained denoiser for purification (zhang2025ddad). Building on this direction, a more principled approach is emerging from a geometric perspective, grounded in the manifold hypothesis (fefferman2016testing). This view posits that adversarial examples are points perturbed off the underlying data manifold, and defense can be achieved by projecting them back (minh2022textual; yang2024adversarial). Recent findings support this, showing that off-manifold examples are critical for robustness, while on-manifold ones aid generalization (altinisik2024explaining). Our work is directly inspired by this geometric intuition, aiming to develop a formal manifold-based purification framework for NLP, where such methods are still nascent.
@@
A key component of our framework is the detection of adversarial samples. We frame this as an out-of-distribution (OOD) detection problem, where clean embeddings are in-distribution and adversarial ones are OOD. Normalizing Flows (NFs) (rezende2015variational), and their continuous counterparts (CNFs) (chen2018neural), are powerful generative models that learn the probability density of data by transforming it to a simple base distribution (e.g., a Gaussian) via an invertible function. They have been successfully used for OOD detection in various domains by assigning low likelihoods to OOD samples (nalisnick2019deep). However, standard CNFs often struggle with high-dimensional data and complex manifold structures. Recent advancements have addressed this limitation in two key ways. First, by applying flows to a model’s semantically rich feature space rather than to raw inputs, OOD detection performance is significantly improved (cook2024feature). This directly supports our approach of operating on sentence embeddings. Second, and more fundamentally, Riemannian Normalizing Flows equip the latent space with a data-dependent Riemannian metric, allowing the model to better capture the underlying geometry of the data manifold. This leads to more accurate density estimation and more reliable OOD detection (diepeveen2024score).
@@
Our defense strategy is predicated on the hypothesis that adversarial perturbations displace text embeddings from a ”clean” data manifold into distinct, separable regions of the embedding space. This section presents a rigorous empirical investigation to validate this hypothesis. We analyze the geometric and distributional properties of encoder’s embeddings from a BERT-base model  (devlin2019bert) fine-tuned on the SST-2 dataset  (socher2013recursive), comparing clean test set examples against their adversarial counterparts generated by TextFooler (jin2020bert).
@@
Figure 3: Density distribution of Local Intrinsic Dimensionality (LID) values. The distribution for attacked embeddings is visibly shifted towards higher LID values compared to clean embeddings.
@@
To first gain a visual intuition, we project the high-dimensional embeddings into a 2D space using three different techniques: Principal Component Analysis (PCA), t-Distributed Stochastic Neighbor Embedding (t-SNE)  (maaten2008visualizing), and Uniform Manifold Approximation and Projection (UMAP)  (mcinnes2018umap). As shown in Figure 2, all three methods reveal a clear separation between the clean (blue circles) and attacked (red crosses) embeddings. While PCA, a linear projection method, shows a noticeable distributional shift, the non-linear, manifold-aware methods t-SNE and UMAP illustrate a more striking separation. In the UMAP projection (Figure 2(c)), the clean and attacked points form well-defined, dense clusters with minimal overlap. This strong visual evidence suggests that clean and adversarial embeddings occupy distinct regions in the latent space, consistent with our manifold separation hypothesis.
@@
To move beyond visual intuition, we quantify the dissimilarity between the clean and adversarial embedding distributions using three standard metrics: Maximum Mean Discrepancy (MMD) with a Gaussian kernel  (gretton2012kernel), Jensen-Shannon Divergence (JSD) (lin2002divergence), and the Wasserstein Distance. Table 1 reports these distances. For a baseline, we first measure the distance between two random halves of the clean embeddings (”Clean In-Distribution”), which is expectedly low. In stark contrast, the distances between the full set of clean and attacked embeddings (”All Clean vs. All Attacked”) are substantially larger across all three metrics. This pattern holds true when examining each class individually. These results provide strong quantitative evidence that adversarial attacks induce a statistically significant distributional shift, confirming that $P_{clean}$ and $P_{adv}$ are indeed different distributions.
Table 2: Mean LID values and Welch’s t-test results. The significantly higher mean LID for attacked embeddings and the extremely low p-value suggest that adversarial perturbations move points to regions of higher geometric complexity.
Comparison Mean LID (Clean) Mean LID (Attacked) p-value
@@
Finally, we investigate if the geometric structure of the manifolds themselves differs. We use Local Intrinsic Dimensionality (LID) (houle2017local), which estimates the dimensionality of the data manifold in the local neighborhood of a data point. A higher LID suggests greater local complexity.
Figure 3 shows the density distributions of LID values for both clean and attacked embeddings. A clear rightward shift is visible for the attacked distribution. This observation is statistically validated in Table 2. The mean LID of attacked embeddings (28.20) is significantly higher than that of clean embeddings (23.74). A Welch’s t-test confirms this difference is highly statistically significant, with a p-value approaching zero ($p\approx 10^{-43}$). This implies that adversarial perturbations do not just move embeddings to a different location, but systematically push them into regions of the embedding space that are geometrically more complex.
@@
Based on this comprehensive empirical evidence and recent advances in the study of embeddings from language models  (li2025unraveling; mamou2020emergence), we establish two foundational hypotheses that motivate our proposed method:
@@
1:Require: Clean data embeddings $Z_{clean}$, attacked data embeddings $Z_{adv}$, text encoder $f_{\theta}$, downstream classifier $C$, hyperparameters $\lambda_{topo},\lambda_{causal}$.
2: Initialize SR-CNF model parameters $\Theta_{flow}$ and MoE metric parameters $\Theta_{metric}$.
@@
8:  Compute $\mathcal{L}_{topo}$ on $Z_{b}$ and $f^{-1}(Z_{b})$.
9:  Compute $\mathcal{L}_{causal}$ between $C(Z_{b})$ and $C(Z_{corr})$.
10:  Compute total loss $\mathcal{L}_{total}=\mathcal{L}_{NLL}+\lambda_{topo}\mathcal{L}_{topo}+\lambda_{causal}\mathcal{L}_{causal}$.
11:  Update $\Theta_{flow}$ and $\Theta_{metric}$ via gradient descent.
12:end for
13: Determine likelihood threshold $\tau$ on a validation set.
14:// Inference Phase
15:for each input embedding $z_{in}$ to be classified do
16:  Calculate $\ell=\log p(z_{in})$ using SR-CNF.
17:  if $\ell<\tau$ then
18:   $z_{out}\leftarrow\text{Geodesic Purification Solver}(z_{in})$
19:  else
20:   $z_{out}\leftarrow z_{in}$
21:  end if
22:  Obtain final prediction from $C(z_{out})$.
23:end for
@@
Instead of assuming a fixed geometry, we learn a Riemannian metric tensor $G(z)$, a positive definite matrix that defines the local geometry at each point $z\in\mathcal{Z}$. To capture the stratified structure observed in our preliminary study, we parameterize $G(z)$ using a Mixture-of-Experts (MoE) network (jacobs1991adaptive; fedus2022switch). This network consists of:
@@
$\frac{dz(t)}{dt}=v(z(t),t),\quad z(t_{0})\in\mathcal{N}(0,I),$ (2)
@@
The total log-likelihood of a data point $z_{in}$ is then obtained by integrating this quantity along the trajectory from the data point at $t=t_{1}$ back to its latent representation $z_{0}$ at $t=t_{0}$:
@@
Table 3: Main results on adversarial robustness evaluation. We report accuracy on the clean test set (Clean%) and accuracy under attack (Aua%), along with the average number of queries per successful attack (#Query). Higher is better for all metrics. Our method, $MC^{2}F$, achieves the best robustness across all datasets and attacks while maintaining high clean accuracy.
@@
MC2F 95.26 48.50 586.4 54.00 214.3 63.20 112.4
@@
subject to the boundary conditions $\gamma(0)=z_{adv}$ and $\gamma(1)=z_{corr}\in\mathcal{M}_{clean}$. The optimal path $\gamma^{*}$ is the geodesic, and its endpoint $\gamma^{*}(1)$ is our purified embedding $z_{corr}$. This optimization problem is solved using an iterative numerical solver that updates the path to minimize its energy while ensuring the endpoint remains on the learned manifold.
@@
To ensure the learned flow $f$ preserves the global shape of the data manifold and avoids topological tearing, we introduce a loss based on differentiable persistent homology. We compute the persistence diagrams (PD)—topological summaries of data—for a batch of clean embeddings $Z_{clean}$ and their latent counterparts $f^{-1}(Z_{clean})$. The loss is the Wasserstein distance between these diagrams:
$\mathcal{L}_{topo}=W_{p}(\text{PD}(Z_{clean}),\text{PD}(f^{-1}(Z_{clean}))).$ (8)
@@
The purification process $z_{adv}\to z_{corr}$ must be semantically consistent. We frame this from a causal perspective, where purification is an intervention to remove the confounding effect of the adversarial perturbation. To enforce this, we use a loss based on information geometry  (volpi2021changing; kim2022fisher). For a clean embedding $z_{clean}$ and its corresponding purified version $z_{corr}$ (obtained by first attacking $z_{clean}$ and then purifying it), we pass both through a fixed, pre-trained downstream classifier $C(\cdot)$ to get two softmax probability distributions, $p_{clean}$ and $p_{corr}$. The causal loss is the Fisher-Rao distance between them:
@@
We conduct experiments on three widely-used text classification benchmarks:
  • SST-2  (socher2013recursive): The Stanford Sentiment Treebank, a binary sentiment classification task with short sentences.
  • AGNews  (jacovi2018understanding): A four-class topic classification dataset, testing performance on multi-class problems.
  • YELP  (asghar2016yelp): The Yelp polarity dataset, a binary sentiment analysis task with longer, more complex reviews.
@@
To ensure a thorough evaluation, we subject the models to three popular and potent word-level adversarial attacks, implemented using the TextAttack toolkit (morris2020textattack):
  • BERT-Attack (li2020bert): A powerful attack that uses a masked language model (BERT) to generate high-quality, contextually appropriate adversarial substitutions.
  • TextFooler (jin2020bert): A greedy-based attack that identifies and replaces the most salient words with synonyms until the model’s prediction changes.
  • TextBugger (li2018textbugger): A versatile attack that combines word-level substitutions with character-level perturbations (e.g., swaps, deletions) to create adversarial examples.
@@
  • Fine-tune  (devlin2019bert): A standard BERT-base model fine-tuned for each downstream task, representing a non-robust baseline.
  • FreeLB (zhu2020freelb): An enhanced adversarial training method that adds adversarial perturbations to word embeddings within a bounded region to improve generalization and robustness.
  • WLRE (huang2022word): A defense framework that enhances robustness by introducing its own perturbations to the input during training, forcing the model to learn more invariant features.
  • SD (Subspace Defense) (zheng2024subspace): A state-of-the-art defense that projects feature representations onto a learned ”clean” subspace to discard adversarial perturbations.

### v2

@@ Introduction
The remarkable success of Pre-trained Language Models (PLMs) like BERT across numerous Natural Language Processing (NLP) tasks (Devlin et al. 2019) is shadowed by their pronounced vulnerability to adversarial attacks (Jin et al. 2020; Alzantot et al. 2018). These attacks introduce minor, often semantically imperceptible, perturbations to input text, yet can trigger catastrophic prediction errors in the model, which becomes especially critical in text classification (TC). (Vázquez-Hernández et al. 2024; Goyal et al. 2023; Li et al. 2020; Wang et al. 2021) In response, a wide array of defense mechanisms have been developed to enhance model robustness (Gao et al. 2023; Zheng et al. 2024; Asl et al. 2024; Zhu et al. 2019). However, their success is often marred by a critical challenge: improving resilience against adversarial inputs frequently comes at the cost of a significant drop in performance on clean, unperturbed data (Ning and Chai 2024; Wang et al. 2021). This robustness-accuracy challenge not only poses a significant barrier to the reliable deployment of PLMs in safety-critical applications but also highlights a fundamental gap in our understanding of their internal workings.
@@
Figure 1: An overview of our proposed Manifold-Correcting Causal Flow ($MC^{2}F$) framework. The system first employs a Stratified Riemannian Continuous Normalizing Flow (SR-CNF) to model the manifold of clean data (colored points). It then identifies out-of-distribution (OOD) or adversarial samples (red points) that lie off the manifold. The Purification Solver module subsequently corrects these OOD embeddings by projecting them back onto the clean manifold along the shortest path (geodesic), a process visualized by the black arrows. In-distribution embeddings are passed through without modification.
@@
To overcome this challenge, we move beyond heuristic defenses and investigate the geometric properties of the model’s embedding space. We begin with a critical empirical observation: in the high-dimensional embedding space of a BERT-style encoder, representations of clean text are geometrically separable from their adversarial counterparts generated by prevalent attacks (e.g., TextFooler, BERT-Attack) (Subhash et al. 2023; Reif et al. 2019). This suggests that adversarial perturbations induce a measurable geometric and distributional shift, rather than creating indistinguishable overlaps. Consequently, the defense problem can be reframed: instead of a brute-force training task, it becomes a more nuanced geometric challenge of identifying and correcting embeddings that have been perturbed off the ”clean” data distribution.
@@
Building on this insight and inspired by the manifold hypothesis—which posits that high-dimensional data lies on a lower-dimensional manifold—we introduce the Manifold-Correcting Causal Flow ($MC^{2}F$). It is a two-module defense system that operates directly on sentence embeddings, embodying a ”detect-and-correct” method to purify adversarial inputs  (Li et al. 2023; Moraffah et al. 2024).
@@
Our work is positioned at the intersection of adversarial defense in NLP and the geometric analysis of embedding spaces. We review the relevant literature in these areas.
@@
The vulnerability of PLMs has been extensively demonstrated by a variety of adversarial attack methods. These methods typically generate perturbations by replacing, inserting, or deleting characters or words under certain constraints to maintain semantic similarity and fluency. Prominent word-level attack algorithms include TextFooler (Jin et al. 2020), BAE (Garg and Ramakrishnan 2020), and TextBugger (Li et al. 2018), which greedily replace words with synonyms that maximize the victim model’s prediction error. More sophisticated methods, such as BERT-Attack (Li et al. 2020), leverage the victim model itself to find more effective substitutes. These attacks serve as standard benchmarks for evaluating the robustness of defense systems, including our own.
@@
The dominant paradigm, adversarial training (AT), augments data with adversarial examples (Zhu et al. 2019; Madry et al. 2017) yet suffers from high computational costs and accuracy trade-offs (Ning and Chai 2024; Wang et al. 2021). Despite improvements, an “illusion of robustness” often persists due to gradient masking (Gao et al. 2023; Wang et al. 2024; Raina et al. 2024). Alternatively, purification methods have evolved toward continuous embedding denoising (Zheng et al. 2024; Asl et al. 2024), with recent works like DAD balancing clean and robust accuracy via MMD-based detection (Zhang et al. 2025). A complementary geometric perspective relies on the manifold hypothesis (Fefferman et al. 2016), projecting off-manifold adversarial points back to the data manifold (Minh and Tuan 2022; Yang et al. 2024). Motivated by evidence that off-manifold samples drive robustness while on-manifold data supports generalization (Altinisik et al. 2024; Dang and Zhu 2025), we pursue a formal manifold-based NLP purification framework.
@@
Adversarial detection can be cast as an out-of-distribution (OOD) task, where clean and adversarial samples represent in- and out-distribution data, respectively. Generative models like Normalizing Flows (NFs) (Rezende and Mohamed 2015) and Continuous NFs (CNFs) (Chen et al. 2018) excel at this by assigning low likelihoods to OOD samples (Nalisnick et al. 2018). To overcome limitations in modeling complex manifolds, modern approaches enhance performance by applying flows in feature spaces (Cook et al. 2024) or integrating Riemannian metrics for precise geometric density estimation (Diepeveen et al. 2024).
Figure 2: Geometric and distributional analysis of clean and attacked SST-2 test set embeddings. (a-c) 2D visualizations using PCA, t-SNE, and UMAP show that adversarial embeddings form clusters visibly distinct from clean embeddings. (d) The density distribution of Local Intrinsic Dimensionality (LID) values shows that attacked embeddings are visibly shifted towards higher LID values. CL-0/1: Clean samples of class 0/1; AT-0/1: Attacked samples of class 0/1.
@@
Our defense strategy is predicated on the hypothesis that adversarial perturbations displace text embeddings from a ”clean” data manifold into distinct, separable regions of the embedding space. This section presents a rigorous empirical investigation to validate this hypothesis. We analyze the geometric and distributional properties of encoder’s embeddings from a BERT-base model  (Devlin et al. 2019) fine-tuned on the SST-2 dataset  (Socher et al. 2013), comparing clean test set examples against their adversarial counterparts generated by TextFooler (Jin et al. 2020).
@@
To first gain a visual intuition, we project the high-dimensional embeddings into a 2D space using three different techniques: Principal Component Analysis (PCA), t-Distributed Stochastic Neighbor Embedding (t-SNE) (Maaten and Hinton 2008), and Uniform Manifold Approximation and Projection (UMAP) (McInnes et al. 2018). As shown in Figure 2(a-c), all three methods reveal a clear separation between the clean (blue circles) and attacked (red crosses) embeddings. While PCA, a linear projection method, shows a noticeable distributional shift, the non-linear, manifold-aware methods t-SNE and UMAP illustrate a more striking separation. In the UMAP projection (Figure 2(c)), the clean and attacked points form well-defined, dense clusters with minimal overlap. This strong visual evidence suggests that clean and adversarial embeddings occupy distinct regions in the latent space, consistent with our manifold separation hypothesis.
@@
To move beyond visual intuition, we quantify the dissimilarity between the clean and adversarial embedding distributions using three standard metrics: Maximum Mean Discrepancy (MMD) with a Gaussian kernel  (Gretton et al. 2012), Jensen-Shannon Divergence (JSD) (Lin 2002), and the Wasserstein Distance. Table 1 reports these distances. For a baseline, we first measure the distance between two random halves of the clean embeddings (”Clean In-Distribution”), which is expectedly low. In stark contrast, the distances between the full set of clean and attacked embeddings (”All Clean vs. All Attacked”) are substantially larger across all three metrics. This pattern holds true when examining each class individually. These results provide strong quantitative evidence that adversarial attacks induce a statistically significant distributional shift, confirming that $P_{clean}$ and $P_{adv}$ are indeed different distributions.
Comparison MLID (CL) MLID (AT) p-value
@@
Table 2: Mean LID (MLID) values and Welch’s t-test results. The significantly higher mean LID for attacked embeddings and the extremely low p-value suggest that adversarial perturbations move points to regions of higher geometric complexity. CL denotes clean embeddings and AT denotes attacked embeddings.
@@
Finally, we investigate if the geometric structure of the manifolds themselves differs. We use Local Intrinsic Dimensionality (LID) (Houle 2017), which estimates the dimensionality of the data manifold in the local neighborhood of a data point. A higher LID suggests greater local complexity.
Figure 2(d) shows the density distributions of LID values for both clean and attacked embeddings. A clear rightward shift is visible for the attacked distribution. This observation is statistically validated in Table 2. The mean LID of attacked embeddings (28.20) is significantly higher than that of clean embeddings (23.74). A Welch’s t-test confirms this difference is highly statistically significant, with a p-value approaching zero ($p\approx 10^{-43}$). This implies that adversarial perturbations do not just move embeddings to a different location, but systematically push them into regions of the embedding space that are geometrically more complex.
@@
Based on this comprehensive empirical evidence and recent advances in the study of embeddings from language models  (Li and Sarwate 2025; Mamou et al. 2020), we establish two foundational hypotheses that motivate our proposed method:
@@
1:Require: Clean data embeddings $Z_{clean}$, downstream classifier $C$, hyperparameters $\lambda_{topo},\lambda_{causal}$.
2: Initialize parameters $\Theta_{flow}$ for the SR-CNF map $f$ and $\Theta_{metric}$ for the MoE metric $G(z)$.
@@
8:  // Compute latent counterparts by inverting the flow
9:  Compute $\mathcal{L}_{topo}$ on $Z_{b}$ and $f^{-1}(Z_{b})$ (Eq. 8).
10:  Compute $\mathcal{L}_{causal}$ between $C(Z_{b})$ and $C(Z_{corr})$.
11:  Compute total loss $\mathcal{L}_{total}=\mathcal{L}_{NLL}+\lambda_{topo}\mathcal{L}_{topo}+\lambda_{causal}\mathcal{L}_{causal}$.
12:  Update $\Theta_{flow}$ and $\Theta_{metric}$ via gradient descent.
13:end for
14: Determine likelihood threshold $\tau$ on a validation set.
15:// Inference Phase
16:for each input embedding $z_{in}$ to be classified do
17:  Calculate $\ell=\log p(z_{in})$ using SR-CNF.
18:  if $\ell<\tau$ then
19:   $z_{out}\leftarrow\text{Geodesic Purification Solver}(z_{in})$
20:  else
21:   $z_{out}\leftarrow z_{in}$
22:  end if
23:  Obtain final prediction from $C(z_{out})$.
24:end for
@@
Instead of assuming a fixed geometry, we learn a Riemannian metric tensor $G(z)$, a positive definite matrix that defines the local geometry at each point $z\in\mathcal{Z}$. To capture the stratified structure observed in our preliminary study, we parameterize $G(z)$ using a Mixture-of-Experts (MoE) network (Jacobs et al. 1991; Fedus et al. 2022). This network consists of:
@@
$\frac{dz(t)}{dt}=v(z(t),t),\quad z(t_{0})\sim\mathcal{N}(0,I),$ (2)
@@
The total log-likelihood of a data point $z_{in}$ is then obtained by integrating this quantity along the forward trajectory from its latent representation $z_{0}$ at $t=t_{0}$ to the data point $z_{in}$ at $t=t_{1}$:
@@
$\mathbf{MC^{2}F}$ 95.26 48.50 586.4 54.00 214.3 63.20 112.4
Table 3: Main results on adversarial robustness evaluation. We report accuracy on the clean test set (Clean%) and accuracy under attack (Aua%), along with the average number of queries per successful attack (#Query). Higher is better for all metrics. Our method achieves the best robustness across all datasets and attacks while maintaining high clean accuracy.
@@
subject to the boundary conditions $\gamma(0)=z_{adv}$ and $\gamma(1)=z_{corr}\in\mathcal{M}_{clean}$. The optimal path $\gamma^{*}$ is the geodesic, and its endpoint $\gamma^{*}(1)$ is our purified embedding $z_{corr}$.
This optimization problem is solved iteratively. We discretize the path $\gamma(t)$ and use gradient descent on its waypoints to minimize the energy functional $\mathcal{L}[\gamma]$ (Eq. 5), with gradients computed respecting the learned metric $G(z)$ (Eq.  1). The constraint $\gamma(1)\in\mathcal{M}_{clean}$ is operationalized by enforcing $\log p(z_{corr})\geq\tau$ (where $\tau$ is the detection threshold) via a soft penalty term in the minimization objective.
@@
To ensure the learned flow (SR-CNF) $f$ preserves the global shape of the data manifold and avoids topological tearing, we introduce a loss based on differentiable persistent homology. We compute the persistence diagrams (PD), topological summaries of data, for a batch of clean embeddings $Z_{b}$ and their latent counterparts $f^{-1}(Z_{b})$. The loss is the Wasserstein distance between these diagrams:
$\mathcal{L}_{topo}=W_{p}(\text{PD}(Z_{b}),\text{PD}(f^{-1}(Z_{b}))).$ (8)
@@
The purification process $z_{adv}\to z_{corr}$ must be semantically consistent. We frame this from a causal perspective, where purification is an intervention to remove the confounding effect of the adversarial perturbation. To enforce this, we use a loss based on information geometry  (Volpi et al. 2021; Kim et al. 2022). For a clean embedding $z_{clean}$ and its corresponding purified version $z_{corr}$ (obtained by first attacking $z_{clean}$ and then purifying it), we pass both through a fixed, pre-trained downstream classifier $C(\cdot)$ to get two softmax probability distributions, $p_{clean}$ and $p_{corr}$. The causal loss is the Fisher-Rao distance between them:
@@
We evaluate our method on three benchmark datasets: SST-2 (Socher et al. 2013) for binary sentiment analysis on short text; AGNews (Jacovi et al. 2018) for four-class news topic classification; and YELP (Asghar 2016), a more complex binary sentiment task involving longer reviews.
@@
We assess robustness using the TextAttack toolkit (Morris et al. 2020) with three established attacks: BERT-Attack (Li et al. 2020), which leverages BERT for context-aware substitutions; TextFooler (Jin et al. 2020), a greedy strategy using synonym replacement; and TextBugger (Li et al. 2018), which employs both word- and character-level perturbations.
@@
  • Fine-tune  (Devlin et al. 2019): A standard BERT-base model fine-tuned for each downstream task, representing a non-robust baseline.
  • FreeLB (Zhu et al. 2019): An enhanced adversarial training method that adds adversarial perturbations to word embeddings within a bounded region to improve generalization and robustness.
  • WLRE (Huang et al. 2022): A defense framework that enhances robustness by introducing its own perturbations to the input during training, forcing the model to learn more invariant features.
  • SD (Subspace Defense) (Zheng et al. 2024): A state-of-the-art defense that projects feature representations onto a learned ”clean” subspace to discard adversarial perturbations.